OpenStudy (anonymous):

Any one solving h9p1,h9p2,h9p3,lab 9 help!!!

OpenStudy (qaisersatti):

have u tried it yourself... ?

OpenStudy (anonymous):

ya got some answers of h9p1 .25 ???? ???? 0 ??? 0 0 0

OpenStudy (anonymous):

H9P1 0.25 7.5 7.5 0 HELP 0 0 0 H9P2 L*C L/R 1 HELP Help (1/R)*sqrt(L/C) HELP H9P3 b/(2*m) 1/(sqrt(m*1/k)) 1/(2*pi*sqrt(m*1/k)) (2*m)/(sqrt(m*(1/k)))

OpenStudy (anonymous):

LAB9 C = 126.5e-12 R = 25000 in the Fig2, change only D for 58% (H=2m, and C=1u)

OpenStudy (anonymous):

H9P1 Q5 is -23.5619449019 the count is -L*2*pi*f*sin(2*pi*.f*t) 2*pi*f = w0 -15*2*pi*0.25*sin(2*pi*.25*5) = -23.5619449019

OpenStudy (anonymous):

H9P1 a) 0.250122093132 b) 7.5 c) 7.5 d) 0 e) -23.5619449019 f) 0 g) 0 h) 0 H9P2 a) L*C b) L/R c) 1 d) sqrt(1/(L*C))/(2*pi) e) 1/(2*R*C) f) (1/R)*sqrt(L/C) g) 2*sqrt(L/C) H9P3 a) b/(2*m) b) 1/(sqrt(m*1/k)) c) 1/(2*pi*sqrt(m*1/k)) d) (2*m)/(sqrt(m*(1/k)))